# A Categorical Framework for Quantifying Emergent Effects in Network Topology

### v3

@@ 5 Computing the derived functors
In Section 3 we introduced how the cohomology object $L^{1}\Phi(A)$ or $R^{1}\Phi(A)$ is able to encode the emergence of system $A$ in a non-trivial way. In this section we discuss how it is computed, resulting in Proposition 5.3 that leads to a computational measure of emergence.
@@
This means to find the emergence of a system, we need to solve equation (31) or (33).
Now we discuss our choice of functors $\Phi$. Given a class of functors $\{\Phi_{\delta}\}$, the goal is to find those that could sustain emergent effects, and especially those with stronger potentials for emergent effects. The class of functors $\{\Phi_{\delta}\}$ of interest will often neglect the partial structure of the system, which resembles information processing systems where some connections are neglected, a ubiquitous class in both natural and engineered systems.
@@
We can see that such functor can indeed sustain emergent effect. Let $f:V\rightarrow W$ be a morphism between two representations,
[…]
and suppose $f_{1}:\ker V(a)\oplus(\mathbb{C}/\ker V(a))\rightarrow\ker W(a)\oplus(\mathbb{C}/\ker W(a))$, is constructed in the following way:
[…]
@@
then $\Phi(f_{2}):\ker V(a)\rightarrow W(a)$ will be zero map while $f_{2}$ is not a zero map. As a consequence, $\Phi$ will not preserve colimit and hence sustain emergent effect.
@@
We now show that $\Phi$ is indeed a functor. We need to show that $\Phi(id_{X})=id_{\Phi(X)}$ for every $X$ in the category, and $\Phi(g\circ f)=\Phi(g)\circ\Phi(f)$ for all morphisms $f:X\rightarrow Y$ and $g:Y\rightarrow Z$. The first condition is satisfied, since $\Phi(id_{X})$ consists of identity maps on vector spaces in $\Phi(X)$, and so is $id_{\Phi(X)}$. The second condition is satisfied as well. We can focus on a single node. $\Phi(g\circ f)$ will be the restriction of $g\circ f$ at $\ker a$. For $x\in\ker a$, $\Phi(g\circ f)(x)=g\circ f(x)=g(f(x))$, on the other hand, $\Phi(g)\circ\Phi(f)(x)=\Phi(g)(\Phi(f)(x))=\Phi(g)(f(x))$. Since $\Phi(g)$ is the restriction of $g$ at $\ker b$, we now need to show that $f(x)\in\ker b$. The morphisms by definition has to satisfy the commutative diagram with the arrow map, so for $x\in\ker a$, we must have $b\circ f(x)=0$, which means $f\in\ker b$. Hence $\Phi(g)(f(x))=g(f(x))$, and we conclude $\Phi(g\circ f)=\Phi(g)\circ\Phi(f)$.
@@
Formally, also consider the multiple edges case, the functor we constructed can be defined in the following way: given a set $E$ of edges to be neglected, we consider the functors
@@
  • $\Phi_{l}:\textbf{Rep}(Q)\rightarrow\textbf{Rep}(Q)$, which maps $V(x)$ to $V(x)/\bigcup_{he=x}\text{Im}V(e)$, where $\bigcup_{he=x}\text{Im}V(e)$ is the union of images of linear maps associated to all edges coming into node $x$. $\Phi_{l}$ maps $V(e),e\in E$ to zero maps. $\Phi_{l}$ maps the morphisms between quiver representations $f:V\rightarrow W$ to their restrictions on $\Phi_{l}(V)$, $f|_{\Phi_{l}(V)}:\Phi_{l}(V)\rightarrow\Phi_{l}(W)$.
@@
  • $\Phi_{r}:\textbf{Rep}(Q)\rightarrow\textbf{Rep}(Q)$, which maps $V(x)$ to $\bigcap_{te=x}\ker V(e)$, where $\bigcap_{te=x}\ker V(e)$ is the union of images of linear maps associated to all edges coming from node $x$. $\Phi_{l}$ maps $V(e),e\in E$ to zero maps. $\Phi_{r}$ maps the morphisms between quiver representations $f:V\rightarrow W$ to their restrictions on $\Phi_{r}(V)$, $f|_{\Phi_{l}(V)}:\Phi_{r}(V)\rightarrow\Phi_{r}(W)$.
We can see that $\Phi_{l}$ and $\Phi_{r}$ equivalently deletes edges in $E$ from a quiver representation.
@@
Note that from a modeling perspective, these functors are also related to the mapping from one scale to a higher scale in a real-world system. When doing a higher level partial observation, some information about the lower level system will be missing, for example, some edges and nodes in the network. Such functor might be further modified into one that represents a coarse-graining scheme or a convolutional neural network.
@@
It’s easy to check that $\Phi_{l}$ and $\Phi_{r}$ are additive functors. Proposition 5.25 in Rotman & Rotman (2009) shows a functor $F:\textbf{Mod}\rightarrow\textbf{Ab}$ is left exact when it preserves kernel and right exact when it preserves cokernel. We can see that for a morphism $f:V(x)\rightarrow W(x)$, we must have $\text{Im}f\subset\ker W(a)$. This means when restricting $f$ on $\ker V(a)$, injectiveness will be preserved while surjectiveness might not be preserved; when restricting $f$ on $V(x)/\ker V(a)$, surjectiveness will be preserves while injectiveness might not be preserved. Hence $\Phi_{r}$ preserves cokernel and is right exact, $\Phi_{l}$ preserves kernel and is left exact.
Now we give the following proposition that quantifies emergence of such a system:
@@
The first left derived functor of $\Phi_{r}$ is
$L^{1}\Phi_{r}(V)=\bigoplus_{a\in E}\Phi_{r}(V(ta)\otimes P_{ha})$ (37)
where $E$ is the set of edges being neglected by $\Phi_{r}$, $P_{ha}$ is the path algebra spanned by all paths from the head of the edge $a$.
@@
$R^{1}\Phi_{l}(W)=\bigoplus_{a\in E}\Phi_{l}(W(ha)\otimes I_{ta})$ (38)
where $E$ is the set of edges being neglected by $\Phi_{l}$, $I_{ha}$ is the path algebra spanned by all paths to the tail of edge $a$.
@@
This proposition informs us that under the choice of the functor that deletes some edges in the network, emergence as $L^{1}\Phi(A)$ or $R^{1}\Phi(A)$ encodes information about the nodes and the paths being affected by deleting these edges. We show in the next Section how we can build a computational measure of emergence from this.
@@ 6 A Numerical Measure of Emergence
$\dim L^{1}\Phi_{r}(V)$ $=\dim\bigoplus_{e\in E}\Phi_{r}(V(te)\otimes P_{he})$
$=\sum_{e\in E}\dim\Phi_{r}(V(te))\times\dim\Phi_{r}(P_{he})$ (39)
@@
$\dim R^{1}\Phi_{l}(W)$ $=\dim\bigoplus_{e\in E}\Phi_{l}(W(he)\otimes I_{te})$
$=\sum_{e\in E}\dim\Phi_{l}(W(he))\times\dim\Phi_{l}(I_{te})$ (40)
Here $\dim\Phi_{r}(V(te))$ and $\dim\Phi_{l}(W(he))$ is the dimension of the image of the vector space $V(te)$ and $W(he)$ under the functor, and $\dim\Phi_{r}(P_{he})$ and $\dim\Phi_{l}(I_{te})$ is the dimension of the image of the path algebra $P_{he}$ and $I_{te}$ under the functor.
@@
Let us discuss how this measure can apply to to graph theory. A direct graph can be regarded as a quiver representation where the vector space associated with each node is 1-dimensional. This kind of quiver representation is considered in, for example, (Armenta & Jodoin, 2021; Ganev & Walters, 2022). Now based on equations (39), we can neglect the $\dim\Phi_{r}(V(te))$ term, and the measure of emergence will base only on the path algebra term, $\dim\Phi_{l}(P_{he})$, which is the number of paths from $he$ that does not contain edges in $E$. Since the vector space associated to the vector spaces are 1-dimensional, if a node $x$ is the head of an edge in $E$, then $V(x)$ will be zero in this case. Let $H$ be the node sets of $\Phi(G)$, then $H$ contains the nodes that are not the head of any $e\in E$. The paths in $\Phi_{l}(P_{he})$ will be all paths in $\Phi(G)$ from $he$ to $H$. We now have an edge-based measure of emergence:
@@
where $N_{H}(x)$ is the set of downstream neighbors of $x$ in $H$.
Note that this result assumes the functor $\Phi$ as deleting nodes in $G\setminus H$. The measure of emergence here could have different forms under other choices of $\Phi$, which will be explored in future work. This result quantifies emergence as counting the number of paths generated from the nodes not preserved by the partial observation. It provides us with a numerical measure of emergence that is based on the local network structure, which is the path originating from each node in $G\setminus H$. When we increase the size of network $G$, the number of paths often grows exponentially, which resembles the scaling laws that are present in a lot of systems where emergent behaviors are observed.
@@
Random Boolean network can be regarded as a generalization of cellular automata Gershenson (2004), and has been used as models to study emergent behavior, especially in the context of biological and complex systems(see, Kauffman, 1993; Gershenson, 2003). In Gershenson & Fernández (2012), random Boolean networks of different connectivity has been simulated, and the information loss as in (43) has been proposed as a measure of emergence, where $I_{out}$ is the Shannon information measured at the end state of the network and $I_{in}$ is the Shannon information measured at the initial state of the network. To compare our measure of emergence with this measure, we run our simulation on networks with a range of connectivities $K$ where the edges and activation functions are randomly generated. We measure emergence of such random Boolean networks by regarding the functor $\Phi_{l}$ as the one that sends a fully connected network $G$ to subnetwork $H$ which remains active during the network dynamics. A node is considered inactive here when it comes into state $1$ in less than $5\%$ of the time after the network dynamics reaches steady state. We consider the quiver representation for each network that has dimension $1$ for the input and output space of each node, and as discussed above, and emergence reflects the number of paths being ”impacted” by the functor. We compared our node based measure of emergence with the information-theoretic measure of emergence as information loss in Figure 4.
Figure 4: Comparison of emergence and information loss at different network connectivity $K$.
@@
From the simulation, we can see that our measure of emergence correlates with information loss, as a way to meassure emergence on the random Boolean networks of different connectivity. The difference here is emergence seems to favor $K=2$ over $K=1$, contrary to information loss. We think there is a connection between our measure of emergence and information loss, although it is not proved at the current stage. As an interpretation, the functor $\Phi$ we choose sends the fully connected network $G$ to the subnetwork $H$ of active nodes. Since the active nodes are given by the steady state of network dynamics, the functor ties the steady state of some connectivity $K$ network to the steady state of the fully connected network. Emergence in this context thus captures the difference between the steady states, and more generally the impact connectivity $K$ has on network dynamics.
We plot emergence together with the percentage of inactive nodes in Figure 5. We can observe that emergence largely correlates with the percentage of inactive nodes in a random Boolean network. Although emergence is dependent on the specific positions of the inactive nodes, we can still see a correlation between emergence and the average number of inactive nodes. This agrees with our intuition in Section 2 and equation (42). We can also see that, our measure of emergence suggests that the network has strongest emergence when the connectivity is neither too dense or too sparse (in the example of random Boolean networks, when $K=2$), like a small-world type network latora2001efficient (Latora & Marchiori, 2001).
Figure 5: Emergence and the percentage of inactive nodes at different network connectivity $K$.

### v4

@@ 5 Computing the Derived Functors
In Section 3 we introduced how the cohomology object $R^{1}\Phi(A)$ is able to encode the emergence of system $A$ in a non-trivial way. In this section we discuss how it is computed, resulting in Proposition 5.3 that leads to a computational measure of emergence.
@@
This means to evaluate the emergence of a system, we need to solve equation (31) or (33), which will be given in Proposition 5.3.
Now we discuss our choice of functors $\Phi$. Given a class of functors $\{\Phi_{\delta}\}$, the goal is to find those that could sustain emergent effects, and especially those with stronger potentials for emergent effects. The class of functors $\{\Phi_{\delta}\}$ of interest will often neglect the partial structure of the system, which resembles information processing systems where some connections are neglected, a ubiquitous class in both natural and engineered systems (Mengistu et al., 2016; Uddén et al., 2020).
To construct edge-deleting functors that sustains emergent effect, we build on the following idea: let $e$ be the edge to be neglected, let $te$ be the source node of $e$ and $he$ be the target node of $e$. Then the functor maps $W(te)$ to its subspace $\ker W(e)$ or the functor maps $W(he)$ to its subspace $W(he)/\text{Im}W(e)$, and then maps $W(e)$ to zero map. Additionally, since mapping the vector space to its subspace may cause other arrow maps to loose domain/image, the functor maps those arrow maps to zero. For $e^{\prime}$ such that its domain or image overlaps with $W(te)/\ker W(e)$ or with $\text{Im}W(e)$, $W(e^{\prime})$ is mapped to zero map.
We can see that such functor can indeed sustain emergent effect. Let $V,W$ both be $\mathbb{C}\rightarrow\mathbb{C}$, $f:V\rightarrow W$ takes the following form:
@@
where $a$ is the edge in the quiver, and suppose $f_{1}:\ker V(a)\oplus(\mathbb{C}/\ker V(a))\rightarrow\ker W(a)\oplus(\mathbb{C}/\ker W(a))$, is constructed in the following way:
@@
where $i_{1}$ is the injection map from $\mathbb{C}/\ker V(a)$ to $\mathbb{C}/\ker W(a)$, suppose $\mathbb{C}/\ker V(a)\subset\mathbb{C}/\ker W(a)$. Then $\Phi(f_{1}):\ker V(a)\rightarrow\ker W(a)$ will be zero map while $f_{1}$ is not a zero map. As a consequence, $\Phi$ will not preserve cokernel, hence not right exact (see, e.g., Rotman & Rotman, 2009), sustaining emergent effect (see, Proposition 3.8).
@@
We now show that $\Phi$ is indeed a functor. We need to show that $\Phi(id_{X})=id_{\Phi(X)}$ for every $X$ in the category, and $\Phi(g\circ f)=\Phi(g)\circ\Phi(f)$ for all morphisms $f:X\rightarrow Y$ and $g:Y\rightarrow Z$. The first condition is satisfied, since $\Phi(id_{X})$ consists of identity maps on vector spaces in $\Phi(X)$, and so is $id_{\Phi(X)}$. The second condition is also satisfied. Consider the example below, focusing on a single node, let $f:\mathbb{C}\rightarrow\mathbb{C}$ and $g:\mathbb{C}\rightarrow\mathbb{C}$, defined on $V(ta)$ and $W(ta)$ respectively. As $\Phi$ maps $V(ta)$ to $\ker V(a)$, $\Phi(g\circ f)$ will be the restriction of $g\circ f$ at $\ker V(a)$. For $x\in\ker V(a)$, $\Phi(g\circ f)(x)=g\circ f(x)=g(f(x))$, on the other hand, $\Phi(g)\circ\Phi(f)(x)=\Phi(g)(\Phi(f)(x))=\Phi(g)(f(x))$. Since $\Phi(g)$ is the restriction of $g$ at $\ker W(a)$, we now need to show that $f(x)\in\ker W(a)$. The morphisms by definition has to satisfy the commutative diagram with the arrow map, so for $x\in\ker V(a)$, we must have $W(a)\circ f(x)=h\circ V(a)=0$, which means $f(x)\in\ker W(a)$. Hence $\Phi(g)(f(x))=g(f(x))$, and we conclude $\Phi(g\circ f)=\Phi(g)\circ\Phi(f)$.
${\mathbb{C}}$${\mathbb{C}}$${\mathbb{C}}$${\mathbb{C}}$${\mathbb{C}}$${\mathbb{C}}$$\scriptstyle{f}$$\scriptstyle{V(a)}$$\scriptstyle{W(a)}$$\scriptstyle{g}$$\scriptstyle{h}$ (36)
@@
Formally, also consider the multiple edges case, the functor we constructed can be defined in the following way: given a set $E$ of edges to be neglected, we consider the functors:
@@
  • $\Phi_{l}:\textbf{Rep}(Q)\rightarrow\textbf{Rep}(Q)$, which maps $V(x)$ to $V(x)/\bigcup_{he=x,e\in E}\text{Im}V(e)$, where $\bigcup_{he=x,e\in E}\text{Im}V(e)$ is the union of images of linear maps associated to all edges in $E$ with head $x$ (coming into $x$). $\Phi_{l}$ maps $V(e),e\in E$ to zero maps. For any edge $e^{\prime}$ such that its domain or image overlaps with $\bigcup_{he=x,e\in E}\text{Im}V(e)$, $V(e^{\prime})$ is mapped to zero map. For morphisms, $\Phi_{l}$ maps the morphisms between quiver representations $f:V\rightarrow W$ to their restrictions on $\Phi_{l}(V)$, $f|_{\Phi_{l}(V)}:\Phi_{l}(V)\rightarrow\Phi_{l}(W)$.
@@
  • $\Phi_{r}:\textbf{Rep}(Q)\rightarrow\textbf{Rep}(Q)$, which maps $V(x)$ to $\bigcap_{te=x,e\in E}\ker V(e)$, the intersection of images of linear maps associated to all edges in $E$ with tail $x$ (coming from $x$). $\Phi_{l}$ maps $V(e),e\in E$ to zero maps. For any edge $e^{\prime}$ such that its domain or image overlaps with $V(x)/\bigcap_{te=x,e\in E}\ker V(e)$, $V(e^{\prime})$ is mapped to zero map. For morphisms, $\Phi_{r}$ maps the morphisms between quiver representations $f:V\rightarrow W$ to their restrictions on $\Phi_{r}(V)$, $f|_{\Phi_{l}(V)}:\Phi_{r}(V)\rightarrow\Phi_{r}(W)$.
We can see that $\Phi_{l}$ and $\Phi_{r}$ equivalently delete edges in $E$ and potentially some other affected edges from a quiver representation. We use $E^{\prime}$ to represent all edges in $V$ whose arrow maps are mapped to zero by $\Phi_{l}$ or $\Phi_{r}$. In practice, the quiver representation can be constructed to eliminate overlaping domains/images, such that $E=E^{\prime}$. In the following we work on the category of quiver representation $\textbf{Rep}(Q)$ where we assume $Q$ is a finite quiver with non-trivial edge set, and the functor $\Phi_{l}$ and $\Phi_{r}$ exists.
@@
Note that from a modeling perspective, these functors are also related to the mapping from one scale to a higher scale in a real-world system. When doing a higher-level partial observation, some information about the lower-level system will be missing, for example, some edges and nodes in the network. Such functor might be further modified into one that represents a coarse-graining scheme or a convolutional architecture.
@@
It’s easy to check that $\Phi_{l}$ and $\Phi_{r}$ are additive functors. Proposition 5.25 in Rotman & Rotman (2009) shows a functor $F:\textbf{Mod}\rightarrow\textbf{Ab}$ is left exact when it preserves kernel and right exact when it preserves cokernel. We can see that for a morphism $f:V(x)\rightarrow W(x)$, we must have $\text{Im}f\subset\ker W(a)$, where $a$ is an edge being removed by the functor, with $ta=x$. This means when restricting $f$ to $\ker V(a)\rightarrow\ker W(a)$, injectiveness will be preserved while surjectiveness might not be preserved; when restricting $f$ to $V(x)/\ker V(a)\rightarrow W(x)/\ker W(a)$, surjectiveness will be preserved while injectiveness might not be preserved. Hence $\Phi_{r}$ preserves cokernel and is right exact, $\Phi_{l}$ preserves kernel and is left exact.
Now we give the following proposition that quantifies emergence as such algebraic structure:
@@
Given a set $E$ of edges to be neglected, the first left derived functor of $\Phi_{r}$ is
$L^{1}\Phi_{r}(V)=\bigoplus_{a\in E^{\prime}_{r}}\Phi_{r}(V(ta)\otimes P_{ha})$ (37)
where $E^{\prime}_{r}$ is the set of edges whose arrow maps in $V$ are mapped to zero by $\Phi_{r}$, $P_{ha}$ is the path algebra spanned by all paths from the head of the edge $a$.
@@
$R^{1}\Phi_{l}(W)=\bigoplus_{a\in E^{\prime}_{l}}\Phi_{l}(W(ha)\otimes I_{ta})$ (38)
where $E^{\prime}_{l}$ is the set of edges whose arrow maps in $W$ are mapped to zero by $\Phi_{l}$, $I_{ha}$ is the path algebra spanned by all paths to the tail of edge $a$.
@@
This proposition shows under the choice of the functor that deletes some edges in the network, emergence as $L^{1}\Phi(A)$ or $R^{1}\Phi(A)$ encodes information about the nodes and the paths being affected by deleting these edges. In the next section, we show how we can build a computational measure of emergence from such structures.
@@ 6 A Numerical Measure of Emergence
$\dim L^{1}\Phi_{r}(V)$ $=\dim\bigoplus_{e\in E^{\prime}_{r}}\Phi_{r}(V(te)\otimes P_{he})$
$=\sum_{e\in E^{\prime}_{r}}\dim\Phi_{r}(V(te))\times\dim\Phi_{r}(P_{he})$ (39)
@@
$\dim R^{1}\Phi_{l}(W)$ $=\dim\bigoplus_{e\in E^{\prime}_{l}}\Phi_{l}(W(he)\otimes I_{te})$
$=\sum_{e\in E^{\prime}_{l}}\dim\Phi_{l}(W(he))\times\dim\Phi_{l}(I_{te})$ (40)
Here $\dim\Phi_{r}(V(te))$ and $\dim\Phi_{l}(W(he))$ is the dimension of the image of the vector space $V(te)$ and $W(he)$ under the functor, and $\dim\Phi_{r}(P_{he})$ and $\dim\Phi_{l}(I_{te})$ is the dimension of the image of the path algebra $P_{he}$ and $I_{te}$ under the functor. As discussed in the previous section, the quiver presentation can be constructed to eliminate overlaping domains or images, such that $E^{\prime}_{r}=E^{\prime}_{l}=E$, so hereafter we use $E$ to represent the set of edges being effectively neglected by $\Phi_{r}$ or $\Phi_{l}$.
@@
Now we discuss how this measure can be applied to networks. A direct graph can be regarded as a quiver representation where the vector space associated with each node is 1-dimensional. This kind of quiver representation is considered in, for example, (Armenta & Jodoin, 2021; Ganev & Walters, 2022). Now based on equations (39), we can neglect the $\dim\Phi_{r}(V(te))$ term, and the measure of emergence will base only on the path algebra term, $\dim\Phi_{l}(P_{he})$, which is the number of paths from $he$ that does not contain edges in $E$. Since the vector space associated to the vector spaces are 1-dimensional, if a node $x$ is the head of an edge in $E$, then $V(x)$ will be zero in this case. Let $H$ be the node sets of $\Phi(G)$, then $H$ contains the nodes that are not the head of any $e\in E$. The paths in $\Phi_{l}(P_{he})$ will be all paths in $\Phi(G)$ from $he$ to nodes in $H$. We now have an edge-based measure of emergence:
@@
where $N_{H}(x)$ is the set of downstream neighbors of $x$ in $H$, and the paths only include nodes in $H$. Note that this result assumes the functor $\Phi$ as deleting nodes in $G\setminus H$. The measure of emergence here could have different forms under other choices of $\Phi$, which will be explored in future work.
This result quantifies emergence, a kind of structural nonlinearity as given in Definition 2.1, as counting the number of paths generated from the nodes not preserved by the partial observation. It provides us with a numerical measure of emergence that is based on the local network structure, which is the path originating from each node in $G\setminus H$. When we increase the size of network $G$, the number of paths often grows exponentially, which resembles the scaling laws that are present in a lot of systems where emergent behaviors are observed.
@@
Random Boolean network can be regarded as a generalization of cellular automata Gershenson (2004), and has been used as models to study emergent behavior, especially in the context of biological and complex systems(see, Kauffman, 1993; Gershenson, 2003). In Gershenson & Fernández (2012), random Boolean networks of different connectivity has been simulated, and the information loss as in (43) has been proposed as a measure of emergence, where $I_{out}$ is the Shannon information measured at the end state of the network and $I_{in}$ is the Shannon information measured at the initial state of the network. To compare our measure of emergence with this measure, we run our simulation on networks with a range of connectivities $K$ where the edges and activation functions are randomly generated. We measure emergence of such random Boolean networks by regarding the functor $\Phi_{l}$ as the one that sends a fully connected network $G$ to subnetwork $H$ which remains active during the network dynamics. A node is considered inactive here when it comes into state $1$ in less than $5\%$ of the time after the network dynamics reaches steady state. We consider the quiver representation for each network that has dimension $1$ for the input and output space of each node, and as discussed above, and emergence reflects the number of paths being ”impacted” by the functor. We compared our node based measure of emergence with the information-theoretic measure of emergence as information loss in Figure 4(a).
(a) Comparison of emergence and information loss at different network connectivity $K$.
(b) Emergence and the percentage of inactive nodes at different network connectivity $K$.
Figure 4: Numerical results showing the correlation of emergence and information loss
@@
From the simulation, we can see that our measure of emergence correlates with information loss, as a way to measure emergence on the random Boolean networks of different connectivity. We think there is a connection between our measure of emergence and information loss, although it is not proved at the current stage. As an interpretation, the functor $\Phi$ we choose sends the fully connected network $G$ to the subnetwork $H$ of active nodes. Since the active nodes are given by the steady state of network dynamics, the functor ties the steady state of some connectivity $K$ network to the steady state of the fully connected network. Emergence in this context thus captures the difference between the steady states, and more generally the impact connectivity $K$ has on network dynamics.
We plot emergence together with the percentage of inactive nodes in Figure 4(b). We can observe that emergence largely correlates with the percentage of inactive nodes in a random Boolean network. Although emergence is dependent on the specific positions of the inactive nodes, we can still see a correlation between emergence and the average number of inactive nodes. This agrees with our intuition in Section 2 and equation (42).